OpenStudy (anonymous):

Please help! z1=-3+4i z2=2-1 Perform the indicated operations and write the answer in the form a + bi. z1+z2= ?

OpenStudy (anonymous):

is z2=2-1 or is it z2=2-i ???

OpenStudy (anonymous):

z2=2-i, I'm sorry :/

OpenStudy (anonymous):

z1 + z2 = (-3 + 4i) + (2 - i) = -3 + 4i + 2 - i = -3 + 2 + 4i - i = (-3 + 2) + (4i - i) = -1 + 3i

OpenStudy (anonymous):

it's like adding "like terms" when you add/subtract complex numbers...:)

OpenStudy (anonymous):

Thanks so much!